What would the rate constant of the reaction be given the following set of

experiments?

A + B → C + D

Experiment [A] (M) [B] (M) Initial Rate (M/s)

1 0.040 0.040 2.2x10-4

2 0.040 0.080 4.4x10-4

3 0.080 0.080 8.8 x 10

-4


1
Expert's answer
2021-11-17T07:57:02-0500

Interpreting the values in the tables and aligning them



"\begin{matrix}
 Exp&A(m)&B(m) & initial\>rate\\
 1&0.040&0.040&2.2×10^{-4}\\
2&0.040&0.080&4.4×10^{-4}\\
3&0.080&0.080&8.8×10^{-4}
\end{matrix}"


Working


"\frac{Rate\,2}{Rate\>1}=\frac{4.4×10^{-4}}{2.2×10^{-4}}=\frac{k(0.04)^x(0.08)^y}{k(0.04)^x(0.04)^y}"


"2=2^y\quad\implies y=1"


"\frac{Rate\>3}{Rate\>2}=\frac{8.8×10^{-4}}{4.4×10^{-4}}=\frac{k(0.080)^x(0.080)^y}{k(0.040)^x(0.080)^y}"


"2=2^x\quad \implies x=1"


Substituting data for trial 1,


"2.2\u00d710^{-4}=k(0.04)^1(0.04)^1"


"K=0.1375\>M^{-1}s^{-1}"